find molecular mass of a gas of the specific heats of the gas are cp=0.2 and cv=0.15.

Grade:11

1 Answers

Arun
6 years ago
Dear Naman

As we know.
Cp – Cv = n* R
0.2 – 0.15 = n * 0.0821
0.05/0.0821 = n
n = 0.61 moles
Regards
Arun (askIITiams forum expert)